4x + 4/2 = 3(x + 1) – 2
Answer by funmath(2873)

4x+4/2=3(x+1)-2
4x+2=3(x)+3(1)-2
4x+2=3x+3-2
4x+2=3x+1
-3x+4x+2=-3x+3x+1
x+2=1
x+2-2=1-2
x=-1
Check by substituting -1 in for x in the original equation.
4(-1)+4/2=3(-1+1)-2
-4+2=3(0)-2
-2=0-2
-2=-2 Fantastic!!! We're right!!!!
x=-1 IS the answer!